OpenStudy (anonymous):

y intercept of 2x-4y=-8 ??

OpenStudy (anonymous):

hint: y intercept is the point where the line 2x-4y = -8 crosses or intercepts the y-axis. What is the immediate thing that jumps out at you when you think of the y-axis? thw value of x is 0 on the y-axis

OpenStudy (anonymous):

2(0)-4y=-8 2-4y=-8 -4y=-8-2 -4y/-4=-10/-4 =2.5 ??

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

Nope. How come 2 times 0 is 2 in your equation?

OpenStudy (anonymous):

ohhh ! thanks!

OpenStudy (anonymous):

you are welcome.